OpenStudy (anonymous):

The square root of a quantity is equal to 3 times the area of a circle with a radius of 2 inches. What is the quantity?

OpenStudy (anonymous):

well do u know the fomula for a circle area

OpenStudy (anonymous):

area of circle = pi (radius) ^2 =3.141592625 (2) ^2 =12.566337061 square root x=3(12.566337061) square root x=37.69911184 (square root x=37.69911184) ^2 x=1421.223034